R-GROUPS AND ELLIPTIC REPRESENTATIONS

FOR SIMILITUDE GROUPS

David Goldberg

Abstract. The tempered spectrum of the similitude groups of non-degenerate symplectic, hermitian, or split orthogonal forms defined over p-adic groups of characteristic zero is studied. The components of representations induced from discrete series of proper parabolic subgroups are classified in terms of R-groups. Multiplicity one is proved. The tempered elliptic spectrum is identified, and the relation between elliptic characters appearing in a given induced representation is determined. Those irreducible tempered representations which are not elliptic and not fully induced from elliptic tempered representations are described.

Typeset by AMS-TEX